# Frontiers in Probabilistic Inference: Sampling Meets Learning

### v1

@@ Appendix C Experimental details and additional results
C.1 Ornstein-Uhlenbeck process
${\mathrm{d}}X_{t}=-\alpha X_{t}{\mathrm{d}}t+\sigma{\mathrm{d}}B_{t},$ (35)
with starting state $x_{0}=e_{1}$, termination state $x_{1}=e_{2}$, termination time $T=1$, and diffusion coefficient $\sigma=0.1$. Ornstein-Uhlenbeck bridges are a standard setting in the literature for quantitatively evaluating performance [23, 2, 45], as the ground-truth control $u^{*}(t,x)$ can be calculated in closed-form as
$u^{*}(t,x)=\frac{2\alpha e^{-\alpha(1-t)}}{(1-e^{-2\alpha t})\sigma^{2}}(x_{1}-xe^{-\alpha t}).$ (36)
The training curves shown in Figure 3(a) are for $\alpha=2.0$ and $d=2$. The results in Figure 3(b) comparing the effect of dimension use $\alpha=2.0$, and the results comparing the effect of OU parameter in Figure 3(c) use $d=2$. To enable a fair comparison between the methods, we use the same conservative neural architecture for both, and use the same learning rate, number of gradient update steps, and number of trajectory simulations per step. We used 500 time-discretisation steps, simulated 64 trajectories at each step, and used a learning rate of 1e-4. For Figures 3(b) and 3(c), we ran the algorithms for 10,000 steps; we sometimes found the value for NGDB to start increasing again later in training, so the reported results are the lowest value for NGDB over the 10,000 training iterations. In all experiments, our method trained significantly faster than NGDB, which can be attributed to not requiring backpropagation through the trajectories.
Both methods learn the true drift to a very high degree of accuracy, and we remark that the difference in KL that is visible between the two methods is predominantly due to the behaviour towards the terminal time $T$. The methods behave comparably earlier in time, but the error of NGDB significantly increases approaching the endpoint. We hypothesise this is a consequence of the differing types of training objective; NGDB uses a KL-based objective, while our approach fits to the desired endpoint behaviour directly in the loss construction so is more accurate towards the terminal time. We found that the methods SDB and FB did not learn the bridges successfully in these experiments, which is unsurprising given that they involve rare events under the base dynamics.
C.2 Cox–Ingersoll–Ross model
We now consider computing bridges for the Cox-Ingersoll-Ross (CIR) model [9], which is used in mathematical finance to model the evolution of interest rates. The CIR model [9] is a 1-dimensional stochastic process that evolves according to the SDE
We consider this example as it allows us to verify our method on an SDE with a spatially-dependent diffusion coefficient $\sigma(t,x)=\varepsilon\sqrt{x}$, and also because the ground-truth is again known in closed form, as the transition densities can be written as $ce^{-u-v}\Big(\frac{v}{u}\Big)^{q/2}I_{q}(2\sqrt{uv}),$ where $c=\tfrac{2a}{(1-e^{-a(T-t)})\sigma^{2}},q=\tfrac{2ab}{\sigma^{2}}-1,u=cxe^{-a(T-t)},v=cx_{T}$, and $I_{q}$ is the modified Bessel function of the first kind of order $q$. This again allows us to report ${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$ and verify that the learned diffusion bridges are correct.
We find that taking $\tilde{b}=0$ performs slightly better than $\tilde{b}=-b$, which in turn performs comparably with NGDB. The methods SDB and FB learn reasonably well in this example as the terminal state is well-covered by the unconditioned dynamics, though they are still outperformed by the methods that learn from controlled dynamics.
C.3 Double-well potential
Next, we consider overdamped Langevin dynamics ${\mathrm{d}}X_{t}=-\nabla U(X_{t}){\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}$} using a classical double-well potential $U(x)=v(x^{2}-1)^{2},$ and aim to learn the bridge from $x_{0}=1$ to $x_{1}=-1$; this bridge problem was previously considered in [35]. In light of the discussion in Appendix B.4, this presents a potentially more troublesome problem setting for our method, as the concave peak between the two metastable states can lead to large Jacobian terms. We again report ${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$ relative to the ground-truth, which in 1 dimension can be approximated by numerically solving the backwards Kolmogorov equation.
For each method, we used 500 time-discretisation steps and 4,000 training steps. For our method and NGDB we used a learning rate of 1e-3, and simulated 64 trajectories at each step, and we use the default hyperparameters for SDB. We note that we were unable to obtain good results for the double-well experiments using the FB method; we hypothesise that is due to the behaviour of the adjoint process used for the training samples, which [2] remark can exacerbate the difficulty of learning the correct dynamics for rare events.
We report results in Table 1 for a well height of $v=3.0$, which results in a strong peak between the states. In Figure 6, we also compare the methods for different heights $v$ of the well barrier. We see that the methods that learn from controlled dynamics (that is, NGDB and our method) outperform SDB which learns from the unconditional dynamics; in particular, the performance of SDB becomes significantly worse for the largest value, which is unsurprising as the uncontrolled simulations will rarely cross between the wells in this case. For lower well heights, our method performs comparably with NGDB for both choices of $\tilde{b}$. As the well height increases, we see that taking $\tilde{b}=-b$ performs better; this reflects the discussion in Appendix B.4 and demonstrates how avoiding large growths in the Jacobian terms can improve performance.
C.4 Cell Diffusion Process
Figure 7: Rare event example for cell diffusion process, from [45], showing $X_{t,1}$ ($\bullet$) and $X_{t,2}$ ($\bullet$). Left: NGDB; Right: Ours.
Figure 8: Multi-modal example for cell diffusion process, from [45], showing $X_{t,1}$ ($\bullet$) and $X_{t,2}$ ($\bullet$). Left: NGDB; Right: Ours
We now consider an example in which the reference drift $b$ is not conservative, so following the discussion in Section 3 we include the reference drift in our parameterisation. We consider the cell differentiation model of [44], which has previously been used to provide a qualitative evaluation of diffusion bridge methods in [23, 2, 45]. The model dynamics are given by the 2-dimensional SDE
We consider the three settings used in [45]: a normal event, a rare event, and a multi-modal example. For the normal event, the starting state is $x_{0}=[0.1,-0.1]$, the termination state is $x_{T}=[2.0,-0.1]$, and the termination time is $T=4$. For the rare event, the starting state is again $x_{0}=[0.1,-0.1]$, the termination state is $x_{T}=[1.0,-0.1]$, and the termination time is again $T=4$. For the multi-modal example, the starting state is $x_{0}=[-1.0,-1.0]$, the termination state is $x_{T}=[-1.0,-1.0]$, and the termination time is $T=5$. In all settings, the diffusion coefficient is $\sigma=0.1$, and we train for 1000 steps. For our approach and NGDB, we use a learning rate of 1e-3, and otherwise use the default parameters for the alternative methods.
There is no ground-truth available in this setting, so to provide a quantitative assessment of performance we report ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$ relative to the base process. All methods performed similarly well in the ‘normal’ setting. In the rare and multimodal settings, NGDB appears to perform slightly better than our method, but both performed strongly and significantly outperformed the methods that use unconditioned dynamics. SDB performed reasonably in the rare setting but failed in the multimodal setting, and FB failed to learn the dynamics accurately in both settings. We also provide a qualitative comparison with NGDB in Figures 8 and 8, and observe similar bridge simulations for the two methods.
C.5 Müller-Brown potential
(a) Using $\tilde{b}=0$ gives unstable learning due to the large Jacobian terms, and thus fails to learn the correct dynamics.
(b) Using $\tilde{b}=-b$ avoids large growth in the Jacobians, enabling successful learning of the correct dynamics.
Figure 9: Comparison of the trajectories obtained by taking auxiliary drifts $\tilde{b}=0$ and $\tilde{b}=-b$ in our method, in the Müller-Brown example.
@@ C.5 Müller-Brown potential
${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})\downarrow$ Max Energy

### v2

@@ 2 Background
In the final line, we use a conditional version of the change of measure in (2); see Lemma D.1. Interchanging the derivative and the expectation in (8) is intuitive and well-known in stochastic analysis, and is often referred to as the ‘reparameterisation trick’ in the machine learning literature (Kingma & Welling, 2014). It can be made rigorous using the theory of stochastic flows (Kunita, 1984; Elworthy et al., 1999), by considering the endpoint $X_{T}$ as a deterministic function of $X_{s}$ for a fixed Brownian trajectory $(B_{t})_{t\in[0,T]}$; details are given in Section D.1. We emphasise here that the Jacobian term $J_{T|t}$ is for the base dynamics $\mathbb{P}$, while the expectation is now over the target dynamics $\mathbb{Q}$}.
Unfortunately, for diffusion bridges we cannot make use of the representation in (10) directly, as this expression is not well defined for the singular reweighting $F=\delta_{x_{T}}$. To avoid this singular behaviour at the terminal time $T$, we can ‘invert’ the property at an earlier time $t<T$, using the tower property and the semigroup property of $J_{t|s}$,
$u^{*}(s,x)={\mathbb{E}}_{\mathbb{Q}}[J_{T|s}^{\top}\,\nabla\log F(X_{T})|X_{s}=x]$
$={\mathbb{E}}_{\mathbb{Q}}\big[{\mathbb{E}}_{\mathbb{Q}}[J_{t|s}^{\top}\,J_{T|t}^{\top}\,\nabla\log F(X_{T})|X_{[s,t]}]~|X_{s}=x\big]$
$={\mathbb{E}}_{\mathbb{Q}}\big[J_{t|s}^{\top}\,{\mathbb{E}}_{\mathbb{Q}}[J_{T|t}^{\top}\,\nabla\log F(X_{T})|X_{t}]~|X_{s}=x\big]$
$={\mathbb{E}}_{\mathbb{Q}}\big[J_{t|s}^{\top}\,u^{*}(t,X_{t})|X_{s}=x\big].$
This results in the following self-consistency property of the optimal control $u^{*}$.
Theorem 2.1 (Self-consistency property (I)).
For $0<s<t<T$, the optimal control $u^{*}$ satisfies
$u^{*}(s,x)={\mathbb{E}}[J_{t|s}^{\top}\,u^{*}(t,X^{u^{*}}_{t})|X_{s}=x].$ (SC1)
Intuitively, this property characterises how the control at an earlier time $s$ should relate to its value at later times $t$ along the solution trajectories (see Figure 2). While (10) held only for differentiable $F$, this self-consistency property holds in more general cases, including for the singular choice $F=\delta_{x_{T}}$ corresponding to the diffusion bridge; we show this rigorously in Section D.1.
A Family of Self-Consistency Properties
The property (SC1) is in fact an instance of a wider class of self-consistency properties, which we will detail further in Section 4 and Appendix C. It is these self-consistency properties that will form the basis for our approach. In particular, under settings with more unstable base dynamics, the Jacobian terms $J_{t|s}$ used in (SC1) can become large; we will later consider a second self-consistency property variant (4.1) which provides more stable behaviour in such scenarios.
Combining Self-Consistency with Terminal Conditions
Theorem 2.1 establishes that self-consistency is a necessary property of the optimal control $u^{*}$, but it falls short of sufficiency; indeed, from the derivation we can see it holds for any choice of $F$. For sufficiency, we must combine self-consistency with an additional terminal condition that incorporates information of the solution’s behaviour at the terminal time $T$. In the next section, we will show how this combined self-consistency framework can be used to design methods for learning diffusion bridge dynamics.
Connections to SOC Methods
The self-consistency framework outlined above also provides a class of methods for solving more general Stochastic Optimal Control (SOC) problems. We remark that the recently introduced adjoint matching objective (Domingo-Enrich et al., 2025), which has proven highly effective for fine-tuning flow-based generative models, can be recovered from this self-consistency framework by taking $t=T$ in (SC1) and imposing the terminal condition $u(T,x)=\nabla\log F(x)$. The adjoint matching objective was originally derived through the adjoint equation and removing terms that are zero at the solution; here, we see that it arises naturally from the derivation above. We comment more on connections to adjoint matching and related stochastic optimal control methods in Appendix F.
3 Control Consistency Losses for Diffusion Bridges
In this section, we utilise the self-consistency property (SC1) to introduce a novel methodology for learning diffusion bridges. In light of the discussion above, we must incorporate additional information regarding the known behaviour at terminal time $T$. The following theorem shows that the diffusion bridge control can be characterised as the unique gradient-form control that forces the dynamics to end at $x_{T}$, while satisfying the self-consistency property.
Theorem 3.1.
Consider the controlled dynamics (3) with control drift $u:[0,T]\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$. Suppose that:
  1. (i)
    $u$ satisfies the self-consistency property (SC1).
  2. (ii)
    $u$ is such that it forces the controlled dynamics to terminate at $X_{T}^{u}=x_{T}$;
  3. (iii)
    $u$ is of gradient form for some time $t$, (that is, $u(t,\cdot)=\nabla\phi(\cdot)$} for some scalar function $\phi$).
Then the controlled process $X_{t}^{u}$ is the diffusion bridge.
The proof of the theorem is provided in Appendix D. We remark that the gradient-form condition (iii) is a technical requirement to rule out rotational components in the control drift $u$. The true control drift is of gradient form, however it could be possible for a self-consistent $u$ to include rotational terms. Enforcing the gradient condition at a single time $t$ is sufficient to prevent such terms arising.
Theorem 3.1 provides the theoretical justification for our proposed approach. Namely, we propose to (see Figure 2):
  • optimise for the self-consistency property using a fixed-point style self-consistency objective.
  • enforce the bridging and gradient-form properties by construction.
3.1 Self-Consistency Losses
To encourage satisfying the self-consistency property (SC1) we wish to minimise the expression
$\mathbb{E}\big[$ $\|u(s,X^{u}_{s})-\mathbb{E}[J_{t|s}^{\top}\,u(t,X^{u}_{t})|X^{u}_{s}]\|^{2}\big]$
$=\mathbb{E}\big[\|u(s,X^{u}_{s})-J_{t|s}^{\top}\,u(t,X^{u}_{t})\|^{2}\big]+C,$ (11)
where $C$ is a constant independent of $u$. There are two dependencies on $u$ in (11): through $u$ itself, but also implicitly through the simulated trajectories $X_{t}^{u}$. We thus propose to train using a fixed-point style loss, akin to similar fixed-point losses used in Domingo-Enrich et al. (2025). Namely, we regress the control values $u(s,X^{u}_{s})$ towards the current estimates of the targets $J_{t|s}^{\top}\,\bar{u}(t,X^{\bar{u}}_{t})$, obtained using samples of the current controlled process $X_{t}^{\bar{u}}$; here, $\bar{u}$ denotes the stopgrad operator, so that the regression targets are temporarily fixed during the optimisation. Importantly, the resulting iterative algorithm does not require propagation of the gradients through the simulated trajectories, which provides significant scalability benefits.
Note that the relation (SC1) holds for any $s<t$; one can therefore integrate over the values of $t$ according to a chosen weighting schedule $\alpha_{t}$, which governs how strongly the self-consistency property is enforced relative to different future times. This results in the following family of training objectives,
$\mathcal{L}(\theta)$ $=\mathop{{\mathbb{E}}}\limits_{\begin{subarray}{c}s\sim\mathbf{U}_{[0,T]}\end{subarray}}\left[\left\|u_{\theta}(s,X_{s}^{\bar{\theta}})-\mathcal{U}_{s}\right\|^{2}\right],$ (12)
$\textrm{where~~}\mathcal{U}_{s}$ $:=\tfrac{1}{A_{s}}\int_{s}^{T}\alpha_{t}J_{t|s}^{\top}\,u_{\bar{\theta}}(t,X_{t}^{\bar{\theta}})\,{\mathrm{d}}t.$ (13)
Here, we use the short-hand notation $X_{t}^{\theta}:=X_{t}^{u_{\theta}}$, $\bar{\theta}$ denotes the stopgrad operator as before, $\alpha_{t}$ can be any weighting function (see Section 4 for a discussion), and $A_{s}=\int_{s}^{T}\alpha_{t}\,{\mathrm{d}}t$ is a normalisation factor.
Solving for Training Targets
We need to compute the training targets $\mathcal{U}_{s}$ in (13), using trajectories simulated with the current control $u_{\theta}$. In order to implement this efficiently, we can solve for a reverse SDE backwards in time along the simulated paths, started from the known final control $\mathcal{U}_{T-\delta t}=u(T-\delta t,X_{T-\delta t})$ (see Section 3.2; this is similar to Pidstrigach et al. (2025, Appendix A), Li et al. (2020); Domingo-Enrich et al. (2025)). To do so, we write the target $\mathcal{U}_{s}$ in terms of the neighbouring target $\mathcal{U}_{s+\delta t}$ by utilising the semigroup property of the Jacobian terms $J_{t|s}$ (see Section B.2), giving the recursive formula
$\mathcal{U}_{s}=\tfrac{1}{A_{s}}\int_{s}^{T}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t})\,{\mathrm{d}}t$
$\approx J_{s+\delta t|s}^{\top}\,\big[\tfrac{A_{s+\delta t}}{A_{s}}\mathcal{U}_{s+\delta t}+(\delta t)\tfrac{\alpha_{s+\delta t}}{A_{s}}u(s+\delta t,X_{s+\delta t})\big]$
$=J_{s+\delta t|s}^{\top}\,\big[\lambda_{s}\mathcal{U}_{s+\delta t}+(1-\lambda_{s})u(s+\delta t,X_{s+\delta t})\big]$ (14)
where in the final line we set $\lambda_{s}=\tfrac{A_{s+\delta t}}{A_{s}}$. Each target $\mathcal{U}_{s}$ is constructed as a weighted combination of the ‘running’ next-step target $\mathcal{U}_{s+\delta t}$ and the next-step control $u(s+\delta t,X_{s+\delta t})$}, before transforming with the single-step Jacobian $J_{s+\delta t|s}\approx{\mathrm{Id}}+\nabla b(s,X_{s})(\delta t)+\nabla\sigma(s,X_{s})\cdot\delta B_{s}$} (approximated with a time discretisation of (5)). This procedure can be carried out using vector-Jacobian products, meaning that the training targets $\mathcal{U}_{s}$ can be computed in an efficient and scalable manner without the need for evaluating the Jacobian matrices explicitly.
3.2 Parameterisation of Bridging Processes
In light of Theorem 3.1 we must also enforce the bridging and gradient-form properties from conditions (ii) and (iii). We will do so by construction, through an appropriate choice of neural parameterisation and by leveraging knowledge of the endpoint $x_{T}$. In this section, we consider neural parameterisations of the full drift $f=b+(\sigma\sigma^{\top})u$} of the controlled process; as $\sigma\sigma^{\top}$ is invertible by assumption, we can compute the corresponding control $u_{\theta}$ used in the training objective (12).
Bridging Property
Since the termination point is known, in the final discretisation step we can jump directly to $x_{T}$. By doing so, we directly enforce the bridging property, and the control function $u(T-\delta t,\cdot)$ at the penultimate timestep is completely determined by construction. This provides the required terminal information, which is then propagated backwards by the self-consistency loss (12) to learn the control at the earlier timesteps.
Choice of Neural Parameterisation
There is considerable freedom in the drift parameterisation; the exact form can be chosen by the user and may be application-dependent. A simple and natural choice that we found to perform well in practice is to parameterise the drift as
$f_{\theta}(t,X_{t})=\tfrac{x_{T}-X_{t}}{T-t}+\sigma(t,X_{t})\,\eta_{\theta}(t,X_{t}),$ (15)
where $\eta$ is a neurally-parameterised vector field. The first term is the drift of a Brownian bridge which guides the controlled process towards the terminal point $x_{T}$, and the second is a neural ‘adjustment’ that we learn.
In many applications, the reference drift $b$ is of gradient form and the diffusion term $\sigma$ is independent of $x$. In such cases, jumping to the endpoint at the final step enforces the gradient-form property of $u$ at time $T-\delta t$, satisfying the requirements of Theorem 3.1. We often found it beneficial to enforce the gradient form of $u$ for all $t$ directly, by taking $\eta=\nabla\psi_{\theta}$ for a scalar-valued neural network $\psi$. When the true drift $b$ is not of gradient form, one can instead ensure $u$ is of gradient-form by including the base drift in the parameterisation. One can also make suitable adjustments to cover other cases; we provide an extended discussion of parameterisation choice in Section B.1.
Algorithm 1 Control Consistency Diffusion Bridge (CCDB)
Input: Self-consistency property (e.g. (SC1) or (4.1)), drift parameterisation $f_{\theta}=b+(\sigma\sigma^{\top})u_{\theta}$, number of iterations NN, batch size $B$, weighting schedule $\alpha$.
for $n=0$ to N1N-1 do
  Path simulation: Simulate $B$ paths of $(X_{t}^{\theta})_{t\in[0,T]}$ according to the controlled dynamics (3), using the current control $u_{\theta}$;
  Target computation: For the chosen self-consistency property, compute training targets $\mathcal{U}_{s}$ backwards along the obtained paths (using e.g. (14) or (49));
  Parameter update: Perform gradient step on $\theta$ using the self-consistency loss (12).
end for
Algorithm
The resulting procedure, which we term Control Consistency Diffusion Bridge (CCDB), is outlined in Algorithm 1. We provide additional implementation details in the next section and in Appendix B.
4 A General Class of Self-Consistency Losses
The presentation so far provides a clean and intuitive self-consistency property (SC1) that can be used to learn diffusion bridges, and in Section 5 we will see that the resulting algorithm often performs extremely well in practice. Here, we shed light on potential issues that can arise in certain settings and present strategies to mitigate them, including a generalisation of the self-consistency property.
Behaviour of the Jacobian Process
Recall that the Jacobian process $J_{t|s}$ evolves as
Note that the increments depend on the current value of $J_{t|s}$. In particular, when $\nabla b$ or $\nabla\sigma$ have positive eigenvalues then the value will increase along those directions. Compounding of such changes over time can lead to explosion of the value of $J_{t|s}$ and subsequently in the targets $J_{t|s}^{\top}\,u(t,X_{t}^{u})$, potentially causing numerical issues during training.
To illustrate this effect conceptually, consider overdamped Langevin dynamics ${\mathrm{d}}X_{t}=-\nabla U(X_{t}){\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}$. The process traverses the scalar potential $U$ randomly, with a tendency to move ‘downhill’. In such a setting, positive eigenvalues occur passing over a peak or ridge. Recalling that $J_{t|s}=\nabla_{X_{s}}X_{t}$, this is indeed intuitive—a perturbation of $X_{s}$ within a valley has little effect on the later position $X_{t}$, but a perturbation made while crossing a peak can compound to cause large deviations in the subsequent trajectory.
Role of $\alpha$ Schedule
Recall that the self-consistency loss (12) depends on a choice of weighting function $\alpha$. Intuitively, the function $\alpha$ determines how ‘far into the future’ the algorithm looks when trying to enforce the self-consistency property. The $\alpha$ schedule therefore provides a trade-off; looking far into the future provides better propagation of the known information from the final step at $T-\delta t$, while a shortened schedule can mitigate growth in the training targets, which can help stabilise training in cases where the Jacobian terms would become large. Following Pidstrigach et al. (2025), we consider three types of $\alpha$ schedule, which differ in how far into the future the algorithm attempts to enforce self-consistency:
  • Next-step prediction: Fits to the training target from only the next discretisation step, so $\mathcal{U}_{s}=J_{s+\delta t|s}^{\top}\,u(s+\delta t,X_{s+\delta t}^{u})$. This corresponds to $\lambda=0.0$ in (14).
  • Average prediction: Fits to the average over the subsequent times $t>s$, so $\mathcal{U}_{s}=\tfrac{1}{T-s}\int_{s}^{T}J_{t|s}^{\top}\,u(t,X_{t}^{u}){\mathrm{d}}t$. This corresponds to a constant $\alpha$ schedule.
  • Final-step prediction: Fits to the training target from only the final step, so $\mathcal{U}_{s}=J_{T-\delta t|s}^{\top}\,u(T-\delta t,X_{T-\delta t}^{u})$. This corresponds to $\lambda=1.0$ in (14).
Generally, we found the average schedule to strike a good balance across a variety of experimental settings, so we use this in our experiments unless otherwise stated. We provide an empirical comparison of $\alpha$-schedules in Section E.7.
Generalising the Self-Consistency Property
We now present an alternative mechanism to combat troublesome behaviour in the Jacobian $J_{t|s}$, by deriving a generalisation of our self-consistency property that retains the same theoretical guarantees.
Recall from the derivation in (8) that the Jacobian $J_{t|s}$ arises from taking the derivative through the $\mathbb{P}$-expectation. We can instead consider a change of measure before performing this operation. Let $\tilde{\mathbb{P}}$ be an ‘auxiliary’ measure associated to the process ${\mathrm{d}}\tilde{X}_{t}=(b+\tilde{b})(t,\tilde{X}_{t}){\mathrm{d}}t+\sigma(t,\tilde{X}_{t}){\mathrm{d}}B_{t}$, in which we include an additional drift term $\tilde{b}$. In the derivation from Section 2, we can instead perform the change of measure
where the Radon-Nikodym derivative is given by the Girsanov weights $\tfrac{{\mathrm{d}}\mathbb{P}}{{\mathrm{d}}\tilde{\mathbb{P}}}=\exp(-\int_{s}^{T}\sigma^{-1}\tilde{b}(r,X_{r})^{\top}{\mathrm{d}}B_{r}-\int_{s}^{T}\tfrac{(\sigma\sigma^{\top})^{-1}}{2}\|\tilde{b}(r,X_{r})\|^{2}{\mathrm{d}}r)$ (Øksendal, 2003, Theorem 8.6.8). Continuing the calculation as in Section 2, we can obtain a generalised self-consistency property (40) (given in Appendix) which instead includes Jacobian terms $\tilde{J}_{t|s}$ for the auxiliary process $\tilde{X}$ rather than the base process $X$, while also picking up additional ‘running cost’ terms.
A natural choice of auxiliary process is to take $\tilde{b}=-b$, so that $\tilde{X}$ has no drift term—when the diffusion coefficient is independent of $x$, this will mean $\tilde{J}_{t|s}={\mathrm{Id}}$. In this case, (40) simplifies to the following property, which we find gives a significant improvement in settings where the growth of the base Jacobian $J_{t|s}$ would cause numerical issues.
Theorem 4.1 (Self-consistency property (II)).
For $0<s<t<T$, the control drift $u^{*}$ of the solution satisfies,
$u^{*}(s,x)={\mathbb{E}}\Big[\int_{s}^{t}$ $\nabla b(r,X_{r}^{u^{*}})^{\top}\,u^{*}(r,X_{r}^{u^{*}})\,{\mathrm{d}}r$
$+u^{*}(t,X_{t}^{u^{*}})\,|X_{s}^{u^{*}}=x\Big].$ (SC2)
Importantly, we retain the same theoretical guarantees when using such instances of the generalised self-consistency property (40) as we do in the previous case.
Theorem 4.2.
The result of Theorem 2.1 holds with the generalised self-consistency property (40) in place of (SC1).
A complete derivation of the generalised self-consistency property and corresponding proofs are given in Appendix D. In Appendix F we also discuss connections with recent developments in the stochastic optimal control literature, and remark that similar variations might be useful for wider SOC problems; we leave investigating this to future work.
(a) KL to solution during training.
(b) Effect of dimension, for $\alpha=2$
(c) Effect of OU parameter, $d=2$.